OpenStudy (anonymous):

When solving systems of equations by elimination do you solve for x or y first???

OpenStudy (anonymous):

It really does not matter... It depends on the given equation if the coefficients of one variable are same in both then we try to eliminate it by adding/subtracting the two equations (depending on the signs of that variable in both the equations. If both the variables have different coefficients in the two equations, then we try to equalise the coefficients for one of them.... I'll give a few examples.....

OpenStudy (anonymous):

2x + 3y = 7 -------- (1) 2x - 7y = -11 -------- (2) Here x has 2 as coefficients in both the equations, so we can subtract (2) from (1) and x will be eliminated If we have 3x + 3y = 7 -------- (1) 2x - 3y = -11 -------- (2) Here y has coefficient +3 in (1) and -3 in (2), so if we add (1) and (2) then y will be eliminated

OpenStudy (anonymous):

Thank you so much. Have a nice day!

OpenStudy (anonymous):

If we have 2x + 3y = 7 -------- (1) 5x - 5y = -11 -------- (2) Here we have two choices i) make coefficients of x same ii) make coefficients of y same If we make coefficients of x same, we do as follows: we multiply (1) by 5 to get 10x + 15y = 35 ------(3) we multiply (2) by -2 to get -10x + 10y = 22 ------(4) Now we add (3) and (4) and x is eliminated If we make coefficients of y same, we do as follows: we multiply (1) by 5 to get 10x + 15y = 35 -------(3) we multiply (2) by 3 to get 15x - 15y = -33 -------(4) Now we add (3) and (4) and y is eliminated Hope this makes it clear for you........

OpenStudy (anonymous):

If u r satisfied with the answer, pls click GOOD ANSWER button for me... (: }))